# Dissecting Open Edge Computing Platforms: Ecosystem, Usage, and Security Risks

### v2

@@ V-E Privacy Risks Against End Users
The information leakage to untrustworthy edge nodes. We also observe that end users of edge-assisted content delivery are subject to information leakage to edge nodes. This is because a typical edge node knows not only the end-user IP addresses but also the plaintext content payloads under delivery. As an edge node in OECP can be malicious, such leakage can enable a potential attacker to link an end-user IP address with the content payloads delivered to it, which can provide a side channel for miscreants for targeted censorship and privacy infringement. For instance, a miscreant can run edge nodes and monitor what videos the nearby residents are watching. Furthermore, as learned from quantitative analysis, a typical edge node tend to deliver content to thousands of end-user IPs on a daily basis and these end-user IPs can be as far as thousands of kilometers from the edge node. What is also observed that the daily content delivery volume of a typical edge node could exceed 15GB. These quantitative results further highlight the concerning and practical impact of this information leakage. Besides, we have also observed the leakage of the device type of end-users, which however, is limited to a subset of content delivery modules. When requesting content, some end-user devices will specify the device type by following the protocol requirements of the respective CDN module, which, along with the leaked IP address, can help an attacker locate the end user more precisely and thus incur a non-negligible privacy concern. Among content delivery flows, 461,446 are subject to the leakage of the device type and have revealed 4,367 different device types.
@@
V-F Insufficient Confinement of OECP Tenants
@@
VI Disclosure and Recommendations
Responsible disclosure. Given aforementioned security and privacy risks, We have conducted responsible disclosure to relevant parties including the OECP operators and the CDN services, for which, contact emails of these relevant parties were collected and disclosure emails have been sent. By this writing, we have received full acknowledgement from OneThingCloud and Xingyu CDN for all the reported security/privacy risks. In the replies, both mentioned that they have reproduced the reported security risks and are working to fix them. For the other parties, we have yet to receive any concrete response.
Security recommendations. Below, we discuss the potential security practices to defend against aforementioned security risks.
VII Discussion
@@
The generalizability of security findings. As we pointed out in §V, some of the identified security risks widely exist in open distributed systems, and become evident when untrusted edge nodes become part of CDN services. For example, the low threat reputation of edge node IPs and the privacy risks against end users have also been observed in cryptojacking and residential proxies [26, 27, 28, 29]. Moreover, the risks related to TLS credentials are generalized to all TLS traffic in open distributed systems. Additionally, other security risks are inherent in the concept of OECPs, such as the the insufficient confinement enforced by edge platforms for their tenants.
Data and code release. We have released the source code for all the tools developed to capture and understand edge traffic  [^3] https://chasesecurity.github.io/Open_Edge_Computing_Platforms/. Regarding the dataset in this paper, to protect privacy in the plaintext traffic and domains, it will be available upon request and necessary background vetting.
VIII Related Works
Edge computing frameworks. Edge computing is a paradigm to offload computing tasks, from either the remote cloud or the resource-constrained end devices, to edge devices that are typically deployed closer to the end devices than the cloud, and are equipped with more computing resources when compared to the end devices. To fulfill the potential of edge computing, various edge computing frameworks tailored for specific scenarios have been proposed and evaluated in previous studies [1, 2, 3]. And these studies can be further divided depending on either the types of end devices or the specific computing tasks. Specifically, typical end devices that are considered in edge computing include AR/VR devices [4], mobile phones [2], IoT devices especially cameras [1, 5], drones [3, 33], connected and autonomous vehicles (CAVs) [34, 6], or even nanosatellites [35, 36]. On the other hand, the tasks to be offloaded to edge nodes include video streaming [2], video filtering [37], video analytic tasks (e.g., object detection) [3, 38, 39, 4, 40], general machine learning training/inference tasks [35, 41, 42], and perception tasks [6, 34, 43] for CAVs, etc. Particularly, Motlagh et al. [3] explored the effectiveness of offloading the video analytic task to edge nodes in the scenario of drone-based crowd surveillance. Furthermore, Liu et al. [4] moved the spotlight to objection detection in devices of augmented reality (AR) and proposed multiple techniques to offload objection detection tasks to edge devices, which has achieved a good trade-off between accuracy and latency. Besides, Zhang, et al. [6] proposed EMP, an edge-assisted perception system that allows multiple nearby vehicles to cooperatively share their raw sensor data with an edge server which in turn merges the resulting data and generates a more complete view of the driving environment. However, despite experimental deployments, these proposals have yet to be deployed in production and their real-world and operational characteristics are unclear.
Measurements on closed edge computing platforms. Xu, et al. [7] conducted the first empirical measurement on a public and operational edge platform, and have revealed multiple interesting observations regarding performance and resource management. For instance, the median round-trip time (RTT) between an end device and the nearest edge node is 10.5ms, which is 1.89× (19.8ms) faster than the nearest cloud site. Also, the studied edge platform was found to have been used in various content delivery activities. However, our study differs from [7] in multiple aspects. First of all, the research subjects are different with regards to who owns the edge computing node. In the platform studied in [7], edge nodes are exclusively deployed and operated by the platform operator, while edge nodes of the two OECPs under our study are contributed by third parties. We thus name the platform studied in [7] as closed edge computing platform since all edge nodes are centrally deployed by the platform itself. Secondly, [7] focused more on the quality of service and workload allocation while our study cares more about the security and privacy implications. Furthermore, even with regards to QoS, our observations are distilled from real-world edge traffic and thus more realistic. while the QoS results of [7] are learned through simulating real-world usage of edge computing.
IX Conclusion
@@
Appendix A Edge Nodes
@@
Appendix B Edge-Assisted Content Delivery
B-A CDN Nodes Observed in Edge Traffic
Table XII presents statistical measurements for nodes of each CDN service as observed in edge traffic. Similar to Table II regarding edge nodes, these data points can only serve as a lower-bound estimate for the scale and distribution of CDN nodes, since not all CDN nodes would communicate with nodes under our control.
TABLE XII: CDN nodes observed in edge traffic.
@@ B-A CDN Nodes Observed in Edge Traffic
B-B End Clients of Content Delivery
Table XIII presents the stats of end-client IPs (i.e., recipient IPs) that have received content payloads from edge nodes under our control.
TABLE XIII: The stats of end-client IPs of content delivery as observed in edge traffic.
@@ B-B End Clients of Content Delivery
As extensively profiled in §IV, edge nodes feature a large scale, a wide distribution as well as the extensive participation in content delivery activities. In these content delivery activities, edge nodes distribute content payloads to either counterpart CDN nodes or end users (e.g., mobile devices). Also, as these content delivery activities involve 6 representative CDN services and 16 popular content providers, a deep understanding of them may shed insightful light on the up-to-date and Internet-wide content delivery. Therefore, in this section, we present such an in-depth characterization of edge-assisted content delivery as observed on edge nodes under our control. Our characterization is focused on answering the following three questions: 1) What content has been distributed and delivered? 2) What usage patterns can we observe for edge-assisted content delivery? 3) What quality of service (QoS) does edge-assisted content delivery achieve?
B-C Content Payloads Under Delivery
@@
TABLE XIV: The delivery stats for content of different types.
@@
Content types. Among content delivered in plaintext through our edge nodes, files of 22 different types have been observed, which can be further grouped as the following content types: video streaming files (e.g., .ts/.mp4/.m4s/.m3u8 files, video/quicktime, video/mp4 types), audio streaming files (e.g., .mp3/.m4a files, audio/mpeg types), image files (e.g., .png/.webp files), program files (e.g., .nds/.json/.bspatch files, text/x-diff types), and general compressed files (e.g., .zip files). Table XIV lists the content types observed for content delivery traffic flows that are in plaintext. For each of these content types, we also present the volume of its content delivery traffic as well as the volume of recipient IPs. One thing to note, as the content types for encrypted delivery traffic flows are invisible to us, this is only a lower-bound estimate for the diversity of content under delivery through edge nodes.
@@
Content size. We also measured the distribution of content files over their file size. As many content delivery traffic flows are encrypted, we cannot directly know the size of such content payloads. Instead, given a content delivery flow, we use the outgoing traffic volume to approximate the size of the content under delivery, which is based on the assumption that the content file under delivery accounts for the majority of the respective outgoing traffic. Table XV presents the cumulative distribution of content delivery flows over the size of their outgoing traffic. And we can see that content payloads tend to be delivered in small chunks, e.g., almost 80% content payloads under delivery have a size less than 1MB while 28% have a size less than 100KB. Typically, the smaller the chunk size, the lower the transmission latency is. Such a small-chunk delivery pattern is consistent with the status quo of Internet-wide content delivery especially for video streaming and live streaming [47]. Then, a CDF plot is also provided in Figure 7 to profile the full-range content size distribution.
TABLE XV: The cumulative distribution of content payloads over their size in MB.
@@
(a) An edge node of TipTime on October 31, 2022.
(b) An edge node of OneThingCloud on November 05, 2022.
Figure 8: Examples of the diurnal distribution of edge nodes with regards to the volume of content delivery.
(a) An edge node of TipTime during a Week.
(b) An edge node of OneThingCloud during a Week.
Figure 9: Examples of the weekly distribution of edge nodes with regards to the daily volume of content delivery.
B-D Content Delivery Patterns
Diurnal/weekly distribution of content delivery activities. Figure 8 presents the typical diurnal evolution of content delivery traffic for edge nodes of TipTime and OneThingCloud respectively. Similar to the content-size measurement, the content delivery volume is calculated as the aggregation of outgoing traffic for all content delivery flows during the respective time window. As we can see that the content delivery traffic varies significantly in volume for different hours of a day. Particularly, for a typical TipTime edge node, the content delivery traffic is at a low-magnitude for most hours of a day before 6 PM, while the average content delivery traffic between 6 PM and midnight has quickly jumped to over 2GB per hour. In contrast, for OneThingCloud, the content delivery activities of a typical edge node can have multiple bursts sparsely distributed across the day, as shown in Figure 8(b), which is likely due to the co-existence of multiple CDN modules from different parties. Also, for both OECPs, there is still much space to further improve the utilization rate of the edge nodes. One way to achieve this is to deploy a combination of CDN modules whose diurnal content delivery patterns complement rather than competing with each other. Similarly, we also profiled the weekly patterns of content delivery activities in terms of delivery volume, as shown in Figure 9. In short, the content delivery of a TipTime edge node is stable across days of a week while that of a OneThingCloud node suffers from a higher extent of fluctuation.
Recipient distribution. As discussed above, our edge nodes have delivered content to 1,129,132 distinct recipient IPs. We further profiled the distribution of these recipient IPs over the volume of content that each of them has received, in an attempt to further understand how our edge nodes are used in content delivery. Here, we measure the volume of content delivery through the outgoing traffic volume of these flows. As shown in Table XVI, the content payloads received by most recipient IPs are of a small volume, e.g., 83.85% recipient IPs have received no more than 1MB content from our edge nodes. Such an observation suggests that a typical end-user in edge-assisted content delivery may concurrently request small content chunks of a large file from many CDN nodes. We also surfaced out top recipient IPs with the largest volume of content delivered, as listed in Table XVII. In contrast with the overall pattern, each of the top 10 recipients has over 1GB content delivered. Our further investigation shows that they are all CDN nodes of Bilibili rather than end users.
TABLE XVI: The cumulative distribution of recipient IPs over the size of content delivery smaller than 1MB.
@@ B-D Content Delivery Patterns
TABLE XVII: Top 10 recipient IPs with the largest volume of content received from our edge nodes.
@@
B-E Quality of Service
We then move to profile the quality of content delivery as fulfilled by edge nodes.
TABLE XVIII: The geographical distance between content recipients and the respective edge node.
Location of the Edge Node $\leq 10\text{KM}$ $\leq 50\text{KM}$ $\leq 100\text{KM}$ $\leq 300\text{KM}$
Shanghai 7.94% 18.59% 21.60% 36.71%
Beijing 4.54% 11.58% 15.38% 45.12%
Shenzhen 2.37% 9.55% 29.08% 55.78%
(a) For an edge node deployed in Shanghai.
(b) For an edge node deployed in Shenzhen.
Figure 10: The geographical distribution of content recipients with regards to the respective edge node (masked with ★).
Locality of content recipients. As expected, it is observed that content recipients of an edge node tend to be located nearby the edge node, which has been well illustrated in Figure 10. Specifically, Figure 10(a) presents a heatmap which profiles the geographical distribution of content recipients for an edge node deployed in Shanghai while Figure 10(b) presents one more heatmap for an edge node deployed in Shenzhen. Apparently, when deciding which end users an edge node can serve, the CDN service will prioritize ones that are closer to the edge node, likely in an attempt to achieve a low latency for the content delivery traffic flows. Table XVIII presents numerical stats about the distance between edge nodes and their respective content recipients, which is aligned with the observations from the heatmaps. For instance, 18% content recipients are within the range of 50KM of an edge node deployed in Shanghai, while an edge node located in Shenzhen has over 55% content recipients distributed within a distance of 300KM. Such locality of content recipients suggests edge-assisted content delivery can place the content closer to the end users when compared with traditional CDN servers, which can likely lead to a lower delay for content delivery flows, as further measured and demonstrated as below.
The delay of content delivery flows. We use the round trip time (RTT) to profile the delay between our edge nodes and remote content recipients. Regarding the calculation of RTT, when a TCP flow is initiated by our edge node, RTT is calculated as the time gap between the outgoing SYN packet and the SYN/ACK packet replied by the remote recipient. When the TCP flow is initiated by the remote recipient, the RTT is calculated as the time gap between the outgoing SYN/ACK packet and the ACK packet that is replied from the content recipient. Table XIX presents the statistical results of the RTT in terms of average, median, standard deviation, and 95th percentile. What is also presented is a direct comparison with RTT results reported in previous works for either traditional CDN networks (Akamai and Edgio) or closed edge CDNs. Compared with these counterparts, the delay of edge-assisted content delivery is either comparable or even lower. For instance, when compared with commercial CDN services, edge-assisted content delivery can achieve a lower 95-percentile delay of 83.37ms and 93.75ms for YunFan CDN and Baidu CDN respectively, while it is 103.38ms for Akamai and 221.74ms for Edgio [25]. One thing to note, the RTT results reported in this study are measured against real-world content delivery traffic flows, while it is not the case for that of these previous studies. Specifically, in [25], the delay under measurement is between open recursive DNS servers and CDN servers while in [7], both the clients and the CDN servers are under the control of the researchers along with content delivery traffic simulated.
TABLE XIX: The deplay of content delivery TCP flows along with a comparison with previous works on commercial CDNs and closed edge platforms.
CDN Flows Round Trip Time (RTT) in ms
Average Median STD $95^{th}$
YunFan CDN 554,376 34.57 20.41 67.79 83.37
Baidu CDN 13,925 43.15 33.02 50.99 93.75
A closed Edge CDN[7] for WiFi users - - 10.50 - -
A closed Edge CDN[7] for LTE users - - 34.20 - -
Akamai [25] - 28.75 8.92 - 103.38
Edgio (formerly known as Limelight) [25] - 81.86 43.15 - 221.74
Appendix C Extra Data for Threat Reputation of Edge Node IPs
@@
TABLE XX: The threat reputation of edge node IPs when measured against malicious traffic flows (MTFs) of the same date.
[online] Edge Group w MTFs $\geq 5$ MTFs $\geq 10$ MTFs Median MTFs TipTime $\text{Edge}_{\text{Traffic}}$ 6.15% 3.30% 2.52% 6 $\text{CDN}_{\text{Traffic}}$ 6.03% 3.26% 2.52% 6 OneThing $\text{Edge}_{\text{Traffic}}$ 2.94% 1.77% 1.32% 6 $\text{CDN}_{\text{Traffic}}$ 4.07% 2.25% 1.71% 6 Both $\text{Edge}_{\text{Traffic}}$ 5.13% 2.82% 2.14% 6 $\text{CDN}_{\text{Traffic}}$ 4.69% 2.57% 1.97% 6
@@
Table XX presents the ratio of edge node IPs that have one or more MTFs observed. One thing to note, an MTF is counted for an edge node IP only if it is captured on a date when the IP is serving as an edge node.

### v3

@@ 5.5 Privacy Risks Against End Users
The information leakage to untrustworthy edge nodes. We also observe that end users of edge-assisted content delivery are subject to information leakage to edge nodes. This is because a typical edge node knows not only the end-user IP addresses but also the plaintext content payloads under delivery. As an edge node in OECP can be malicious, such leakage can enable a potential attacker to link an end-user IP address with the content payloads delivered to it, which can provide a side channel for miscreants for targeted censorship and privacy infringement. For instance, a miscreant can run edge nodes and monitor what videos the nearby residents are watching. Furthermore, as learned from quantitative analysis, a typical edge node tend to deliver content to thousands of end-user IPs on a daily basis and these end-user IPs can be as far as thousands of kilometers from the edge node. What is also observed that the daily content delivery volume of a typical edge node could exceed 15GB. These quantitative results further highlight the concerning and practical impact of this information leakage.
@@
5.6 Insufficient Confinement of OECP Tenants
@@
5.7 Responsible Disclosures
Given aforementioned security and privacy risks, We have conducted responsible disclosure to relevant parties including the OECP operators and the CDN services, for which, contact emails of these relevant parties were collected and disclosure emails have been sent. By this writing, we have received full acknowledgement from OneThingCloud and Xingyu CDN for all the reported security/privacy risks. In the replies, both mentioned that they have reproduced the reported security risks and are working to fix them. For the other parties, we have yet to receive any concrete response.
6 Discussion
Security recommendations. Below, we discuss the potential security practices to defend against aforementioned security risks. While defending against privacy risks, we propose EdgeTor, a protocol that features the use of edge nodes as relays to form anonymous circuits towards specific content payloads or edge nodes. Besides, we recommend that fresh threat intelligence should be integrated on edge nodes. Then, short-lived and node-specific credentials can mitigate the exposure of credentials. The details of recommendations can be found in Appendix F.
The generalizability of security findings. As we pointed out in §5, some of the identified security risks widely exist in open distributed systems, and become evident when untrusted edge nodes become part of CDN services. For example, the low threat reputation of edge node IPs and the privacy risks against end users have also been observed in cryptojacking and residential proxies [24, 25, 26, 27]. Moreover, the risks related to TLS credentials are generalized to all TLS traffic in open distributed systems. Additionally, other security risks are inherent in the concept of OECPs, such as the the insufficient confinement enforced by edge platforms for their tenants. More related works can be found in Appendix G.
@@ 6 Discussion
7 Data Availability.
We released the source code for all the tools developed to capture and understand edge traffic in https://chasesecurity.github.io/Open_Edge_Computing_Platforms/. Regarding the dataset in this paper, to protect privacy in the plaintext traffic and domains, it will be available upon request and necessary background vetting.
8 Conclusion
@@
9 Acknowledgments
This work is jointly supported by the National Natural Science Foundation of China under Grant No. 62302473 and No. 62372268, University of Science and Technology of China through the Innovation Fund for Young Investigators, the Shandong Provincial Natural Science Foundation (ZR2021LZH007 and ZR2022LZH013), the Jinan City ”20 New Universities” Funding Project (2021GXRC084), and the Key R&D Program of Shandong Province, China (2024CXGC010114).
Appendix A Collecting and Analyzing Edge Activities
Deployment options to choose. When designing this framework, the first decision we should make is about what node deployment option to choose, this is because both TipTime and OneThingCloud have offered multiple options for deploying their edge nodes, so as to accommodate for as many categories of devices as possible. Specifically, Both of them provide edge node payload as firmware (mainly for router devices), docker images, Android apps while OneThingCloud even further supports running an edge node as a plugin of a network attached device. One thing to note, given observations and findings distilled from docker-based edge node deployment, it is intuitive to wonder whether they are applicable to edge nodes deployed through other options (e.g., firmware), which we have confirmed through small-scaled and short-period edge node deployment using options other than docker images.
Resource requirements. First, varied across OECPs, content delivery is carried out either through TCP or UDP flows. To support UDP-based content delivery, the edge node should at least be configured with an open NAT so that it can serve incoming UDP packets, e.g., via STUN-based NAT traversal. Furthermore, to support TCP-based content delivery, the edge node should be able to listen to a public IP address and accept incoming TCP connections. To satisfy both, we configure each edge container with the host network type, which allows it to have full access to the network of the host, a virtual machine deployed in a public cloud with static public IP addresses attached.
Besides, a runtime monitoring module is further equipped along the docker container of each edge node so as to capture its activities and resource usage. Particularly, given a running edge node, its resource consumption will be snapshotted every minute through the docker stats command. All these runtime logs will be transmitted to a central backup server on a daily base and get cleaned up locally to free up disk space.
Tricks for Manual Analysis. Below we discuss few tricks while we manually analyze the raw network traffic with Wireshark. One is to utilize the correlation between traffic flows especially the correlation between plaintext traffic flows and encrypted ones. For instance, as revealed by our manual study, in edge nodes of TipTime, the CDN module from YunFan will be deployed as the by-default edge task. And the CDN module periodically queries the control server regarding what content to cache and where to pull the respective content payload. Once the respective content payload(e.g., an Android APK file) is cached, a log message will be sent to the logging server of the same CDN. These traffic flows are causally related. Then, since the last traffic flow for logging is in plaintext, it can thus help us to infer the semantics of the first two traffic flows which are encrypted.
Appendix B Edge Nodes
@@
Appendix C Edge-Assisted Content Delivery
TABLE IX: The list of CDN services and content providers.
CDN OECP Content Provider
Wangsu CDN TipTime Toutiao
Bilibili CDN OneThing Cloud Bilibili
Baidu CDN OneThing Cloud Haokan Video, Baidu Cloud
Xunlei CDN OneThing Cloud Xunlei
CDN Nodes Observed in Edge Traffic. Table X presents statistical measurements for nodes of each CDN service as observed in edge traffic. Similar to Table II regarding edge nodes, these data points can only serve as a lower-bound estimate for the scale and distribution of CDN nodes, since not all CDN nodes would communicate with nodes under our control.
TABLE X: CDN nodes observed in edge traffic.
@@ Appendix C Edge-Assisted Content Delivery
End Clients of Content Delivery. Table XI presents the stats of end-client IPs (i.e., recipient IPs) that have received content payloads from edge nodes under our control.
TABLE XI: The stats of end-client IPs of content delivery as observed in edge traffic.
@@
Appendix D Content Payloads Under Delivery
@@
TABLE XII: The delivery stats for content of different types.
@@
Content types. Among content delivered in plaintext through our edge nodes, files of 22 different types have been observed, which can be further grouped as the following content types: video streaming files (e.g., .ts/.mp4/.m4s/.m3u8 files, video/quicktime, video/mp4 types), audio streaming files (e.g., .mp3/.m4a files, audio/mpeg types), image files (e.g., .png/.webp files), program files (e.g., .nds/.json/.bspatch files, text/x-diff types), and general compressed files (e.g., .zip files). Table XII lists the content types observed for content delivery traffic flows that are in plaintext. For each of these content types, we also present the volume of its content delivery traffic as well as the volume of recipient IPs. One thing to note, as the content types for encrypted delivery traffic flows are invisible to us, this is only a lower-bound estimate for the diversity of content under delivery through edge nodes.
@@
Content size. We also measured the distribution of content files over their file size. As many content delivery traffic flows are encrypted, we cannot directly know the size of such content payloads. Instead, given a content delivery flow, we use the outgoing traffic volume to approximate the size of the content under delivery, which is based on the assumption that the content file under delivery accounts for the majority of the respective outgoing traffic. Table XIII presents the cumulative distribution of content delivery flows over the size of their outgoing traffic. And we can see that content payloads tend to be delivered in small chunks, e.g., almost 80% content payloads under delivery have a size less than 1MB while 28% have a size less than 100KB. Typically, the smaller the chunk size, the lower the transmission latency is. Such a small-chunk delivery pattern is consistent with the status quo of Internet-wide content delivery especially for video streaming and live streaming [32]. Then, a CDF plot is also provided in Figure 7 to profile the full-range content size distribution.
TABLE XIII: The cumulative distribution of content payloads over their size in MB.
@@
Recipient distribution. As discussed above, our edge nodes have delivered content to 1,129,132 distinct recipient IPs. We further profiled the distribution of these recipient IPs over the volume of content that each of them has received, in an attempt to further understand how our edge nodes are used in content delivery. Here, we measure the volume of content delivery through the outgoing traffic volume of these flows. As shown in Table XIV, the content payloads received by most recipient IPs are of a small volume, e.g., 83.85% recipient IPs have received no more than 1MB content from our edge nodes. Such an observation suggests that a typical end-user in edge-assisted content delivery may concurrently request small content chunks of a large file from many CDN nodes. We also surfaced out top recipient IPs with the largest volume of content delivered, as listed in Table XV. In contrast with the overall pattern, each of the top 10 recipients has over 1GB content delivered. Our further investigation shows that they are all CDN nodes of Bilibili rather than end users.
TABLE XIV: The cumulative distribution of recipient IPs over the size of content delivery smaller than 1MB.
@@
TABLE XV: Top 10 recipient IPs with the largest volume of content received from our edge nodes.
@@
Appendix E Extra Data for Threat Reputation of Edge Node IPs
@@
TABLE XVI: The threat reputation of edge node IPs when measured against malicious traffic flows (MTFs) of the same date.
[online]
Edge Group w MTFs $\geq 5$ MTFs $\geq 10$ MTFs Median MTFs
TipTime $\text{Edge}_{\text{Traffic}}$ 6.15% 3.30% 2.52% 6
$\text{CDN}_{\text{Traffic}}$ 6.03% 3.26% 2.52% 6
OneThing $\text{Edge}_{\text{Traffic}}$ 2.94% 1.77% 1.32% 6
$\text{CDN}_{\text{Traffic}}$ 4.07% 2.25% 1.71% 6
Both $\text{Edge}_{\text{Traffic}}$ 5.13% 2.82% 2.14% 6
$\text{CDN}_{\text{Traffic}}$ 4.69% 2.57% 1.97% 6
@@
MTFs reports from the proprietary threat intelligence platform. Table XVI presents the ratio of edge node IPs that have one or more MTFs observed. One thing to note, an MTF is counted for an edge node IP only if it is captured on a date when the IP is serving as an edge node.
TABLE XVII: Statistics of edge nodes’ malicious activities as captured by VirusTotal: Mal denotes malicious.
Edge Group Mal W Mal URLs W Malware
TipTime $\text{Edge}_{\text{Traffic}}$ 3.99% 3.96% 0.81%
$\text{CDN}_{\text{Traffic}}$ 3.62% 3.59% 0.73%
$\text{Edge}_{\text{DNS}}$ 2.75% 2.73% 0.91%
OneThing $\text{Edge}_{\text{Traffic}}$ 2.33% 2.33% 0.22%
$\text{CDN}_{\text{Traffic}}$ 3.01% 2.99% 0.46%
$\text{Edge}_{\text{DNS}}$ 3.07% 3.04% 1.31%
Both $\text{Edge}_{\text{Traffic}}$ 3.47% 3.44% 0.62%
$\text{CDN}_{\text{Traffic}}$ 3.21% 3.19% 0.55%
$\text{Edge}_{\text{DNS}}$ 2.78% 2.76% 0.94%
As shown in Table XVII, a non-negligible fraction of edge node IPs are considered as malicious by VirusTotal for either hosting malicious URLs or distributing malware payloads. Particularly, among edge node IPs directly observed in edge traffic ($\text{Edge}_{\text{Traffic}}$), 3.47% are detected as malicious by VirusTotal. We then look into top IPs that are most reported by VirusTotal, and we observe that 49 out of the top 50 IPs used to be bots of the Mozi IoT botnet since they have one or more URLs detected for Mozi payload distribution. For instance, 183.15.89.96, a TipTime edge node observed in traffic, used to distribute Mozi payloads and has 20 URLs detected by VirusTotal, e.g., http://183.15.89.96:38713/Mozi.m, and https://183.15.89.96:38713/mozi.a. Among IPs having malicious URLs detected, over 25% have such kinds of Mozi payload URLs.
Appendix F Security Recommendations.
Appendix G Related Works
Measurements on closed edge computing platforms. Xu, et al. [7] conducted the first empirical measurement on a public and operational edge platform, and have revealed multiple interesting observations regarding performance and resource management. For instance, the median round-trip time (RTT) between an end device and the nearest edge node is 10.5ms, which is 1.89× (19.8ms) faster than the nearest cloud site. Also, the studied edge platform was found to have been used in various content delivery activities. However, our study differs from [7] in multiple aspects. First of all, the research subjects are different with regards to who owns the edge computing node. In the platform studied in [7], edge nodes are exclusively deployed and operated by the platform operator, while edge nodes of the two OECPs under our study are contributed by third parties. We thus name the platform studied in [7] as closed edge computing platform since all edge nodes are centrally deployed by the platform itself. Secondly, [7] focused more on the quality of service and workload allocation while our study cares more about the security and privacy implications.